### v1

@@ Abstract
Effective code optimization in compilers plays a central role in computer and software engineering. While compilers can be made to automatically search the optimization space without the need for user interventions, this is not a standard practice since the search is slow and cumbersome. Here we present CodeZero, an artificial intelligence agent trained extensively on large data to produce effective optimization strategies instantly for each program in a single trial of the agent. To overcome the huge range of possible test programs, we prepare a large dataset of training programs that emphasize quality, naturalness, and diversity. To tackle the vast space of possible optimizations, we adapt deep reinforcement learning to train the agent in a sample-efficient manner through interacting with a world model of the compiler environment. Evaluation on both benchmark suites and production-level code optimization problems demonstrates our agent’s supercompiler performances and zero-shot generalization abilities, outperforming built-in optimization options designed by compiler experts. Our methodology kindles the great potential of artificial intelligence for engineering and paves the way for scaling machine learning techniques in the realm of code optimization.
@@ 1 Introduction
Entering the post-Moore’s law era, code optimization is crucial for computer and software engineering, which plays an important role in realizing the full potential of slow-growing hardware. Developers typically rely on a compiler’s ability to transform input programs into semantically equivalent but more efficient versions, improving metrics like execution time, code size, and power consumption. For example, standard optimization options -O1, -O2, and -O3 aim to reduce execution time, while the -Os and -Oz options are crafted to reduce code size. Still, it is not common for users to explore beyond these conventional compiler options. Given the vast diversity of programs and platforms, coupled with the increasing number of optimization passes integrated into compiler frameworks, these off-the-shelf optimization strategies predefined heuristically by compiler experts may struggle to guarantee near-optimal performance across ever-changing scenarios triantafyllis2003compiler ; georgiou2018less .
Automatic code optimization is therefore crucial in compilers. Autotuning bodin1998iterative improves code by systematically searching the optimization space through iterative executing and profiling optimization strategies. This search technique can yield remarkable performance gains but must be rerun for each new program with thousands of compilations, which is too time-consuming to be practical for all but a few specialized use cases. Meanwhile, machine learning techniques hold the capabilities to generalize the optimization strategy of one program to other similar ones, thereby facilitating faster code optimization. A direct method is to use supervised learning to predict good optimization strategies of input programs calder1997evidence ; fursin2008milepost ; zanella2020yacos , which is impractical due to prohibitive computations to construct labeled training data by search. Another more promising routine, reinforcement learning, that successfully discovered faster sorting algorithm (AlphaDev fawzi2022discovering ) and matrix multiplication algorithm (AlphaTensor mankowitz2023faster ), can explore the optimization space from feedback on optimization metrics without requiring optimal labeled data. For both techniques, broad generalization across different programs, even out of the training samples, arises as a major bottleneck. The community has noted that in a range of machine learning applications devlin2018bert ; brown2020language ; radford2021learning ; kirillov2023segment , training high-capacity models on large-scale datasets has yielded unprecedented performances. For example, large language models like GPT-4 have demonstrated impressive zero-shot generalization abilities ouyang2022training ; openai2023gpt4 . However, in the area of code optimization, the current practice is to learn optimization heuristics in a per-program manner huang2019autophase ; shahzad2022reinforcement or from a small training set with hundreds of programs mammadli2020static ; jain2022poset , lagging far behind the new era of solving challenging problems by scaling up machine learning models.
@@
Figure 1: CodeZero agent performs code optimization using a learned policy in a compiler environment. Guided by a learned policy, at each step, CodeZero analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics. Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
In this study, we focus on the LLVM lattner2004llvm phase ordering problem, a longstanding challenge for compiler research, and propose CodeZero, a reinforcement-learned code optimization agent, capable of generating a sequence of optimization passes tailored to a particular input program. The problem is formulated such that a code optimization agent, upon encountering a program, selects a series of optimization passes, and receives feedback based on the outcomes of applying these transformations, from the compiler environment. To tackle the huge range of possible test programs, we aim to enhance the generalization ability of trained agents by assembling a large-scale training dataset of natural programs. Several previous works rely on randomly generated programs due to data scarcity but the significant distribution shift from random-generated and human-written programs can even hurt the generalization of learned agents when tested on real-world scenarios cummins2022compilergym . Our training data not only includes real-world programs sourced from GitHub da2021anghabench but also incorporates complex algorithmic solutions of competitive programming li2022competition and diverse programs generated by large language models (LLM) tihanyi2023formai . To tackle the vast space of possible optimizations ($\sim 10^{73}$ sequences), we employ a state-of-the-art model-based reinforcement learning method hafner2023mastering to train the code optimization agent sample-efficiently. This method not only learns a predictive world model of compilers to reduce the amount of real compiler executions but also benefits generalization by learning representations that better capture the structure of the compiler state transitions anand2021procedural . After training on massive programs via trial-and-error, the code optimization agent can generalize in a zero-shot way to previously unseen programs with superior optimization performance against off-the-shelf compiler heuristics.
We demonstrate the effectiveness of our trained CodeZero agent on a range of domains cummins2022compilergym from benchmark suites encompassing fundamental algorithms that are ubiquitously employed in everyday applications, to production-level open-source programs, including object files from C++ TensorFlow abadi2016tensorflow and OpenCV culjak2012brief library. On six test datasets, our agent can produce optimization sequences in a single trial, yielding more efficient code size reduction compared to the -Oz flag. Detailed analysis underscores the effectiveness of both the newly introduced program datasets and the model-based reinforcement learning technique within the realm of code optimization. We posit that the agents developed through our approach could be integrated into the existing toolkits of optimization strategy in compilers alongside other manually designed heuristics, such as -Oz or -O3.
2 CodeZero Agent
2.1 Code Optimization as Decision Making
As illustrated in Figure 1, compilers consist of three main components: the front end which translates the source code into an intermediate representation (IR), the middle end, and the back end which converts IR to the binary code. The middle end is responsible for language- and platform-agnostic optimizations over the IR, implemented as passes to either collect information about the program or apply a transform on it, like function inlining, loop unrolling, and dead code elimination. For example, the LLVM-opt tool has more than 100 optimization passes available. A specific order of applying these passes forms an optimization sequence for an input program. This is critical as the right selection and ordering of passes can significantly boost the program’s performance. Despite that compiler developers have provided standard optimization sequences at various levels, e.g. -O2, -O3, -Oz, these preset sequences may not always yield optimal results, especially for emerging programs written in new frameworks such as TensorFlow. Particularly due to the increasing number of optimization passes, it is an open challenge to determine the most effective sequence for each program.
This problem, known as phase ordering, can be naturally formulated as a partially observable Markov decision process (POMDP) sutton2018reinforcement . In this formulation, a code optimization agent determines the optimization sequence of an input program through a series of interactions with the compiler environment, guided by its policy $\pi$. The process starts at the initial state $s_{0}$, representing the IR of the program to be optimized, which is randomly sampled from all IRs of interest. It is critical that aiming to capture the most important characteristics of the target optimization, the agent only receives partially observable information $o_{0}$ of the state. The observation space can vary widely, ranging from manually designed features (e.g. the number of basic blocks) huang2019autophase to more complex tree-based or graph-based program representations cummins2020programl ; alon2018general , and even raw text strings of IR cummins2017end . At each time step $t=0,1,2,\dots$, the agent takes an action $a_{t}$ based on its policy, corresponding to selecting an optimization pass. Following this action, the agent receives an immediate reward $r_{t+1}$, reflecting changes in the optimization metrics, and an observation $o_{t+1}$ of the next state $s_{t+1}=p(s_{t},a_{t})$ which represents the IR transformed by the compiler using the selected pass. This process can be terminated either when the agent finds no positive gains can be achieved or a maximum number of steps is reached. The goal of the agent is to learn an optimization policy $\pi(a_{t}\mid o_{\leq t})$ that effectively maximizes the cumulative rewards, thereby optimizing the performance metrics of the final IR.
2.2 Large-Scale Data Preparation
To ensure that our CodeZero agent can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset. Firstly, the dataset must reflect naturalness. Training data should be within the distribution of human-written programs, otherwise, overfitting to programs that deviate significantly from this, such as those generated by tools like Csmith yang2011finding and llvm-stress lattner2004llvm , could provide no benefits or even hurt the generalization to real-world scenarios. Secondly, diversity in the dataset is essential. It should encompass a wide range of human-written program styles and structures, ensuring globally comprehensive coverage of possible scenarios. Lastly, the pursuit of high-quality training data is imperative. Existing large-scale program collections da2021anghabench ; cummins2020programl have been proposed to serve as training data cummins2022compilergym but are proven unfruitful in our preliminary experiments. Instead, we focus on data that features moderate lengths of IR, complex algorithmic logic, and potential room for optimizations. This allows the agent to flexibly explore and understand the local transformation structure within the IR space. Collectively, these three properties – naturalness, diversity, and high quality – contribute to aligning the visited IR distribution during training with that will be encountered in real-world applications.
We construct our training dataset containing hundreds of thousands of programs by combining three distinct single-source program datasets, CodeContests li2022competition , FormAI tihanyi2023formai , and AnghaBench da2021anghabench . CodeContests comprises human-written solutions to competitive programming problems with complex and optimizable algorithmic logic. FormAI is a large collection of AI-generated C programs with various functionality types and coding styles, aiming to enrich the dataset’s diversity. AnghaBench is a collection of real-world C programs extracted from GitHub. As shown by visualization in Figure 7, our training data has a broad coverage of evaluation programs in a variety of domains. Future expansions of the dataset can be conducted following the aforementioned principles.
2.3 Agent Training with Model-based RL
Figure 2: Design overview of CodeZero. (a) In the model-based training loop, the CodeZero agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations. (b) Deployment of the trained agent is capable of zero-shot generalization to unseen programs, delivering effective optimizations. (c) The compiler environment is set up with the input program’s IR, $s_{0}$. At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements. (d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal states, and predicting future observations and rewards in response to input actions. (e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
@@ 2.3 Agent Training with Model-based RL
While existing work on reinforcement learning for code optimization prevalently focuses on model-free RL methods, model-based RL can offer advantages in terms of both sample efficiency sutton2018reinforcement and generalization anand2021procedural . Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming. This is further compounded when constructing complex observations, such as control-data flow graphs brauckmann2020compiler ; cummins2020programl . Model-based RL addresses these challenges by learning a world model to approximate state transitions and reward signals of the environment. This allows the agent to learn its policy by simulating trajectories based on model predictions, rather than relying solely on trial-and-error interactions in the real compiler environment. This approach thus improves sample efficiency. Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy mazoure2021cross .
We train the CodeZero agent by adapting an advanced model-based RL method, Dreamer hafner2023mastering , as depicted in Figure 2. This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment by approximating the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$. Through imaginary rollouts using this world model over a horizon $H$, the policy $\pi_{\psi}(a_{t}|o_{\leq t})$ can be learned using the REINFORCE policy gradients williams1992simple with an entropy regularizer $\mathrm{H}$. The training objective is formulated as
$\mathcal{L}\left(\psi\right)\doteq\mathbb{E}_{\hat{p}_{\theta},\hat{r}_{\theta},\pi_{\psi}}\left[\sum_{t=1}^{H}-\left(V_{t}-v_{\xi}\left(o_{\leq t}\right)\right)\log\pi_{\psi}\left(a_{t}\mid o_{\leq t}\right)-\eta\mathrm{H}\left[\pi_{\psi}\left(o_{\leq t}\right)\right]\right],$ (1)
where $V_{t}$ is the imagined return estimation and the value prediction $v_{\xi}$ serves as a baseline for variance reduction greensmith2004variance . The value prediction is learned separately by regressing the cumulative rewards that the agent should expect to receive from the current state. All components are realized as deep neural networks.
After extensive training on a wide range of programs, our agent demonstrates the capability to generalize “zero-shot” — that is, to be effective without further training — to new, unseen programs across various domains.
3 Results
3.1 Evaluation Benchmarks
Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems. This focus is driven by the practical advantages of code size as a metric: it is both cost-effective and convenient to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
We evaluate our method on benchmarks from the CompilerGym platform cummins2022compilergym : benchmark suites including cBench fursin2007midatasets , CHStone hara2008chstone , MiBench guthaus2001mibench , and NASA Parallel Benchmarks (NPB) bailey1995parallel , in addition to kernels from open source projects including BLAS lawson1979basic , Linux, OpenCV culjak2012brief , and TensorFlow abadi2016tensorflow . We disregard benchmarks generated by synthetic program generators lattner2004llvm ; yang2011finding since they do not align with real-world scenarios.
Figure 3: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
Table 1: Top performances of the zero-shot CodeZero agent on individual programs. The agent’s optimization sequences are streamlined, omitting passes that do not contribute to performance enhancement. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
Dataset Program Pass Sequences by CodeZero Agent Code size
O0 CodeZero Oz (Reduction)
cBench sha -sroa -gvn -instcombine -simplifycfg -licm -early-cse -simplifycfg -indvars -gvn -simplifycfg -memcpyopt -reassociate 799 349 500 (1.43$\times$)
bzip2 -sroa -gvn -simplifycfg -instcombine -simplifycfg -early-cse -simplifycfg -reassociate -memcpyopt -jump-threading -functionattrs -licm -gvn -simplifycfg -reassociate -early-cse -jump-threading -early-cse -instcombine -simplifycfg 28748 13565 15946 (1.18$\times$)
qsort -sroa -gvn -simplifycfg -instcombine -simplifycfg -early-cse -gvn -indvars -gvn -simplifycfg 638 280 315 (1.12$\times$)
OpenCV #41 -lowerinvoke -simplifycfg -globalopt 28 18 28 (1.56$\times$)
#9 -sroa -gvn -sroa -early-cse -simplifycfg -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -simplifycfg -jump-threading -early-cse -reassociate -instcombine 9510 6341 9269 (1.18$\times$)
TensorFlow #17 -sroa -gvn -simplifycfg -lowerinvoke -simplifycfg -sroa -early-cse -indvars -simplifycfg 5512 4247 5450 (1.28$\times$)
#6 -gvn -simplifycfg -lowerinvoke -simplifycfg -instcombine -early-cse -gvn -jump-threading -instcombine 16173 12643 16076 (1.27$\times$)
3.2 Code Optimization Results
Figure 3 reports CodeZero’s performance in reducing code size, measured in terms of IR instruction count, compared to the expert-designed -Oz flag in LLVM. Remarkably, CodeZero, without any specific training on in-domain programs, outperforms -Oz across all but two test benchmark datasets in a single trial. The marginal performance variations between different methods on the BLAS and Linux datasets suggest that they are already highly optimized. CodeZero also demonstrates the ability to match or even slightly surpass a competitive autotuning method, random search, within a similar time budget (in the order of seconds), while the latter blindly aggregates hundreds of trials. This notable performance, coupled with the significant advantage over a single random trial, underscores the effectiveness of our agent’s policy learning. Moreover, CodeZero’s zero-shot generalization capability either matches or exceeds the performance of in-domain counterparts, which are trained using the train subset of each benchmark dataset. This is particularly significant in cases like the NPB dataset, which contains only 22 training programs. In such scenarios, where data sparsity presents a challenge for in-domain agents, the zero-shot CodeZero notably achieves an extra 3% reduction in code size. Notably, CodeZero’s pre-trained policy can generalize to IRs compiled from novel programming languages beyond C and C++, as evidenced in its performances on the BLAS and NPB datasets of Fortran programs. Further, in an AI-generated benchmark dataset of Objective-C, another language supported by LLVM-Clang apart from C/C++, CodeZero successfully improves upon LLVM -Oz, achieving an average code size reduction of 1.027$\times$, and reaching up to 2.87$\times$ in certain instances.
3.3 Program Case Study
We analyze the internal behavior of our CodeZero agent to optimize IR code via recording the sequence of passes chosen by the policy on individual programs. Table 1 presents the agent’s top performance outcomes on various benchmark datasets. We observe that the agent indeed produces a specialized optimization strategy tailored for each program. Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -gvn (global value numbering), and -simplifycfg, as particularly effective in code size optimization.
3.4 Effect of Training Dataset
Figure 5: Evaluations of CodeZero with different datasets (a-b) and RL algorithms (c-d). (a) Performance of CodeZero trained on various datasets, with “Mixture” representing a combined dataset of CodeContests, FormAI, and AnghaBench. (b) Performance of CodeZero when trained on the full CodeContests dataset versus a subset of 100 randomly selected training programs from CodeContests. (c) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set. A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends. (d) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets.
In our quest to understand the impact of training datasets on generalization, we explored how different datasets from various domains and sizes influence performance. As shown in Figure 5a, we discovered that the CodeContests dataset stands out as the most effective for generalizing to our test benchmarks. Although comparable outcomes were observed with other datasets like FormAI and AnghaBench, agents trained on these datasets notably lag in performance when tested on OpenCV and TensorFlow benchmarks. Moreover, combining these datasets does not yield performance enhancement. Consequently, we present in Figure 3 the results based on the agent trained exclusively with the CodeContests dataset. These results underscore the significance of high-quality training data, a principle that is increasingly recognized as vital in other scalable machine learning applications, such as large language models. In Figure 5b, we compare the agent trained on the full CodeContests dataset versus the one trained on a 100-program subset sampled from the same dataset. This subset is comparable in size to the datasets used in previous research mammadli2020static ; jain2022poset that applies reinforcement learning to code optimization. The suboptimal performance on this smaller dataset highlights its inadequacy for training a strong zero-shot generalizable agent and emphasizes the significance of a larger and more diverse training set prepared by this study.
3.5 Sample Efficiency and Zero-Shot Generalization
We further evaluate the sample efficiency and zero-shot generalization abilities of CodeZero based on the world model algorithm, Dreamer, against a range of model-free counterparts. These include notable algorithms such as DQN mnih2015human , A2C mnih2016asynchronous , APEX horgan2018distributed , IMPALA espeholt2018impala , and PPO schulman2017proximal (see Appendix B.4 for more details). Figure 5c shows that while PPO is the most competitive among the model-free baselines, Dreamer outperforms it by learning an order of magnitude faster, requiring fewer interactions with the compiler environment to achieve excellent performance. Additionally, as Figure 5d indicates, Dreamer excels in generalizing to unseen test benchmarks, outstripping its counterparts. These findings verify our hypothesis that a world model-based agent holds a stronger capability to capture the environment’s structure and can effectively enhance zero-shot generalization.
4 Discussion
We tackled the major challenge of zero-shot generalization that arises when applying artificial intelligence techniques to improve code optimization. We introduce the CodeZero agent that leverages the strengths of a diverse, high-quality training program dataset, as well as the sample efficiency and generalization capabilities of world model-based reinforcement learning. Our results have shown that CodeZero, with its zero-shot generalization ability, attains supercompiler code optimization performance, in the challenging phase ordering problem for code size reduction. Our study provides an AI-centric framework and serves as a meaningful step towards scaling machine learning techniques for code optimization. There is a substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world models, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
5 Methods
In this section, we provide the mathematical and computational details of our approach.
5.1 LLVM Phase Ordering POMDP

### v2

@@ Abstract
Effective code optimization in compilers is crucial for computer and software engineering. The success of these optimizations primarily depends on the selection and ordering of the optimization passes applied to the code. While most compilers rely on a fixed sequence of optimization passes, current methods to find the optimal sequence either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training. We introduce CompilerDream, a model-based reinforcement learning approach to general code optimization. CompilerDream comprises a compiler world model that accurately simulates the intrinsic properties of optimization passes and an agent trained on this model to produce effective optimization strategies. By training on a large-scale program dataset, CompilerDream is equipped to serve as a general code optimizer across various application scenarios and source-code languages. Our extensive experiments first highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard. More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and other state-of-the-art methods in both settings of value prediction and end-to-end code optimization.
@@ 1 Introduction
Code optimization plays an important role in realizing the full potential of software and hardware. Developers desire a universal solution to transform input programs into semantically equivalent but more efficient versions without manual effort. To achieve this, compilers usually employ a front-end that translates the source code of any language into an intermediate representation (IR), a back-end that converts IR to the binary code optimizer, and a middle-end optimizer that performs language- and platform-agnostic optimizations over the IR of the original code (Figure 1). The optimizer is typically implemented as a series of passes applying transforms on the code. The performance of the optimization primarily depends on the selection and order of the optimization passes. Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance. For example, -O1, -O2, and -O3 improve for execution speed, and -Os and -Oz for reducing program size. Obviously, given the vast diversity of programs and platforms, these off-the-shelf optimization strategies predefined by compiler experts are not optimal for most circumstances, and we can significantly improve the performance of a specific program if we can find a better optimization pass sequence compared to that provided by compilers (Triantafyllis et al. 2003; Georgiou et al. 2018).
@@
Figure 1: CompilerDream agent performs code optimization using a learned policy in a compiler environment. Guided by a learned policy, at each step, CompilerDream analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics. Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
Automatically optimizing the pass sequence is thus crucial in improving the compiler optimization performance. To be practical, such an algorithm must produce a satisfactory pass sequence within a reasonable time and handle a wide variety of programs. However, current research often fails to meet these requirements simultaneously. Search-based methods (Bodin et al. 1998) ensure near-optimality but require thousands of compilations for each new program, making it too time-consuming for regular use. In contrast, machine learning methods can either predict the optimization sequences directly or estimate the outcomes of applying an optimization sequence to guide a search, thereby avoiding the time-consuming process of interacting with the compiler to validate the optimization outcome. However, in addition to potentially sacrificing some optimality compared to searching, these learning-based methods face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples.
A range of machine learning applications (Devlin et al. 2019; Brown et al. 2020; Radford et al. 2021; Kirillov et al. 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance. Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al. 2022). However, the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al. 2020; Shahzad et al. 2022) or from relatively small training sets comprising only a few hundred programs (Mammadli, Jannesari, and Wolf 2020; Jain et al. 2022). Additionally, the predominant use of model-free reinforcement learning methods with limited capacity may also hinder generalization.
In this study, we focus on the LLVM (Lattner and Adve 2004) phase ordering problem, a longstanding challenge for compiler research, and propose CompilerDream, a world model-based reinforcement learning approach capable of generating an optimization pass sequence tailored to any particular input program. Our model-based method (Hafner et al. 2023) not only learns an accurate predictive world model of compilers to simulate real compiler executions, whose number can be greatly reduced, but also benefits generalization by learning representations that better capture the structure of the compiler state transitions (Anand et al. 2022). Trained on a carefully curated large-scale dataset of natural programs, the world model accurately captures the dynamics of the compiler optimization process, enabling the optimization agent to generalize to unseen programs and predict superior optimization sequences.
We demonstrate the effectiveness of CompilerDream on a range of domains (Cummins et al. 2022), from benchmark suites encompassing fundamental algorithms to production-level open-source programs, including object files from C++ TensorFlow (Abadi et al. 2016) and OpenCV (Culjak et al. 2012) library. Even without considering generalization, CompilerDream’s strong optimization capabilities place it at the top of the CompilerGym leaderboard for autotuning. For general code optimization, the large-scale trained world models of CompilerDream can accurately predict the outcomes of pass sequences on unseen programs, a setting known as value prediction. Beyond selecting optimization sequences based on value prediction, the agent can also directly produce optimization sequences in a single trial, yielding more efficient code size reduction compared to both the -Oz flag and state-of-the-art methods.
The main contributions of this work are three-fold:
  • We introduce CompilerDream, the first model-based reinforcement learning approach for code optimization.
  • We first leverage a large-scale dataset, CodeContests, to achieve strong generalization on unseen programs.
  • Extensive experiments spanning autotuning, value prediction, and end-to-end code optimization, demonstrate the superior performance of CompilerDream.
2 Related Work
One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order. This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation (Bodin et al. 1998) or autotuning (Datta et al. 2008). However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler strategy. This limitation underscores the importance of integrating machine learning techniques.
Supervised Learning
Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins 2020). The first approach requires an extensive search on each training programs to identify the most effective optimization sequence, which then serves as the data labels. An early example (Calder et al. 1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al. 2008), a practical attempt to integrate machine learning into a production compiler, GCC. It employs models trained on a large dataset of programs distributed over the Internet. The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one (Stephenson et al. 2003; Luk, Hong, and Kim 2009).
Reinforcement Learning
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data (Kulkarni and Cavazos 2012). This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al. 2021), loop transformation (Haj-Ali et al. 2020; Brauckmann, Goens, and Castrillón 2021), and graph partitioning (Mirhoseini et al. 2017). Several works relevant to us have explored the full optimization pipeline, i.e. the LLVM phase ordering problem, including AutoPhase (Haj-Ali et al. 2020), CORL (Mammadli, Jannesari, and Wolf 2020), and POSET-RL (Jain et al. 2022).
Model-based Reinforcement Learning
@@ 2 Related Work
While existing work on reinforcement learning for code optimization predominantly focuses on model-free RL methods, our work pioneers the use of an advanced model-based RL approach. Model-based RL (Sutton and Barto 2018) can offer advantages in terms of both sample efficiency and generalization. Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming. This is further compounded when constructing complex observations, such as control-data flow graphs (Brauckmann et al. 2020; Cummins et al. 2021). Model-based RL addresses these challenges by learning a world model that approximates state transitions and reward signals of the compiler. This is related in spirit to the aforementioned approach of supervised performance models but goes further by also learning a policy through simulating trajectories based on model predictions. This approach thus improves sample efficiency not only by eliminating the necessity for a guided search process during deployment but also by reducing trial-and-error interactions with the real compiler environment during training. Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy (Anand et al. 2022; Mazoure et al. 2022; Ma et al. 2024).
Figure 2: Design overview of CompilerDream. (a) In the model-based training loop, the CompilerDream agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations. (b) The trained agent can be deployed and is capable of zero-shot generalization to unseen programs, delivering effective optimizations. (c) The compiler environment is set up with the input program’s IR, $s_{0}$. At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements. (d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal neural states, and predicting future observations and rewards in response to input actions. (e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
3 Method
In this section, we first present the mathematical formulation of the code optimization problem (Section 3.1). Then we introduce the CompilerDream approach for both compiler world model learning (Section 3.2) and optimization agent learning (Section 3.3). Finally, we elaborate on considerations of curating large-scale training dataset (Section 3.4).
3.1 Phase Ordering Decision Process